OpenStudy (anonymous):

write and equation of a line that passes through points 2,3 and 2,7

OpenStudy (e.mccormick):

Do you know how to find the slope and use the point slope formula after that to find the line?

OpenStudy (anonymous):

slope is y1-y2/x1-x2?

OpenStudy (e.mccormick):

Yes! So use that to find the slope first.

OpenStudy (anonymous):

so the slope is 4?

OpenStudy (e.mccormick):

Nope. This one is a bit of a trick.

OpenStudy (anonymous):

is it -4/0?

OpenStudy (e.mccormick):

Well, you get either 4/0 or -4/0, but the slope can only be one thing, not two. Also, remember the rule about dividing by 0?

OpenStudy (anonymous):

so the slope is 0?

OpenStudy (e.mccormick):

Not quite. 0 is a good number here, but for technical reasons it is not correct. Ever hear of things being undefined?

OpenStudy (anonymous):

oh yeah! now i remeber so because its undefined is there no equation?

OpenStudy (e.mccormick):

Well, there is an equation, but the x value stays the same. When they are like this and the slope is like that, it means it is a vertical line.

OpenStudy (e.mccormick):



OpenStudy (e.mccormick):

Now, if the slope becomes 0, then it is a horozontal line.

OpenStudy (anonymous):

Thank you! and if I have to do the same thing for the points 2,4 and 6,16 is the slope 3? and what do I do after that?

OpenStudy (e.mccormick):

They also call vertical an infinite slope.

OpenStudy (e.mccormick):

OK. You know the slope is 3. Now, seen the \(y-y_1=m(x-x_1)\) form of a line?

OpenStudy (e.mccormick):

You just solved for m.

OpenStudy (anonymous):

Where does the 3 go in that equation?

OpenStudy (e.mccormick):

\[m=\frac{y_1-y_2}{x_1-x_2}\]m is slope.

OpenStudy (e.mccormick):

So this time around, \(m=3\).

OpenStudy (anonymous):

ok so now do I use y=3x+b? and how do I know waht the other variables are?

OpenStudy (e.mccormick):

You need to use the point/slope form. \(y-y_1=m(x-x_1)\) y and x are the variables. Then take a point, \((x_1,y_1)\) from either \((2,4)\) or \((6,16)\) and plug it in.

OpenStudy (e.mccormick):

Then you can distribute the slope and solve for y, which gets you the y intercept form.

OpenStudy (anonymous):

so I shoud be doing y-4=3(x-2) or 16-4=3(6-2)?

OpenStudy (e.mccormick):

The first.

OpenStudy (anonymous):

ok thank you!

OpenStudy (e.mccormick):

What do you get when done with that?

OpenStudy (e.mccormick):

\[y-y_1=m(x-x_1)\implies \\ y=mx-mx_1+ y_1\implies \\ y=mx+b\] You get the same b no matter which point you use. You just use only one point for the last part, not both. \[y-16=3(x-6)\implies \\ y-16=3x-18\implies \\ y=3x-18+16\implies \\ y=3x-2\] or \[y-4=3(x-2)\implies \\ y-4=3x-6\implies \\ y=3x-6+4\implies \\ y=3x-2\] One last tip! Watch out for negative points on the grid. You can't just ignore the negative sign. It has to go in as part of the variable. So if had been say a -2 it would have been \((x-(-2))\) so \((x+2)\).